OpenStudy (anonymous):

how to simplify sqrt 18/4 show me the work the ans is 3 sqrt2/2

OpenStudy (anonymous):

\[\sqrt{\frac{18}{4}}\]you are allowed to apple the root to numerator and denominator separately\[=\frac{\sqrt{18}}{\sqrt{4}}\]the numerator of 18 has a perfect square factor of 9, and the denominator is a perfect square\[=\frac{\sqrt{9*2}}{2}\]appl the radical to the 9 in the radicand\[=\frac{3\sqrt{2}}{2}\]and this is the answer

OpenStudy (anonymous):

THANK YOU